OpenStudy (anonymous):

Simplify the expression. csc^2x sec^2x/sec^2x + csc^2x

OpenStudy (anonymous):

is the equation: csc^2x sec^2x/(sec^2x + csc^2x) or (csc^2x sec^2x/sec^2x) + csc^2x

OpenStudy (anonymous):

The top one

OpenStudy (anonymous):

csc^2x = 1/sin^2x sec^2x = 1/cos^2x sin^2x + cos^2x = 1 what can you do with them?

OpenStudy (anonymous):

\[\frac{ (\csc ^{2}x)(\sec ^{2} x) }{ \csc ^{2}x+\sec ^{2} x }=\frac{ (\frac{ 1 }{ \sin ^{2}x })(\frac{ 1 }{ \cos ^{2}x }) }{ \frac{ 1 }{ \sin ^{2}x }+\frac{ 1 }{ \cos ^{2}x } }=\frac{ 1 }{ \sin ^{2}x \cos ^{2}x(\frac{ 1 }{ \sin ^{2}x }+\frac{ 1 }{ \cos ^{2}x }) }=\frac{ 1 }{ \frac{ \sin ^{2}x \cos ^{2}x }{ \sin ^{2}x } + \frac{ \sin ^{2}x \cos ^{2}x }{ \cos ^{2}x }}\]

OpenStudy (anonymous):

\[\frac{ 1 }{ \frac{ \sin ^{2}x \cos ^{2}x }{ \sin ^{2}x } + \frac{ \sin ^{2}x \cos ^{2}x }{ \cos ^{2}x }}=\frac{ 1 }{ \cos ^{2}x+\sin ^{2}x }=\frac{ 1 }{ 1 }=1\]

OpenStudy (anonymous):

Thank you!